# Quantifying Aleatoric and Epistemic Dynamics Uncertainty via Local Conformal Calibration

### v2

@@ Abstract
Whether learned, simulated, or analytical, approximations of a robot’s dynamics can be inaccurate when encountering novel environments. Many approaches have been proposed to quantify the aleatoric uncertainty of such methods, i.e. uncertainty resulting from stochasticity, however these estimates alone are not enough to properly estimate the uncertainty of a model in a novel environment, where the actual dynamics can change. Such changes can induce epistemic uncertainty, i.e. uncertainty due to a lack of information/data. Accounting for both epistemic and aleatoric dynamics uncertainty in a theoretically-grounded way remains an open problem. We introduce Local Uncertainty Conformal Calibration (LUCCa), a conformal prediction-based approach that calibrates the aleatoric uncertainty estimates provided by dynamics models to generate probabilistically-valid prediction regions of the system’s state. We account for both epistemic and aleatoric uncertainty non-asymptotically, without strong assumptions about the form of the true dynamics or how it changes. The calibration is performed locally in the state-action space, leading to uncertainty estimates that are useful for planning. We validate our method by constructing probabilistically-safe plans for a double-integrator under significant changes in dynamics. Companion website: um-arm-lab.github.io/lucca.
@@ 1 Introduction
This paper proposes Local Uncertainty Conformal Calibration (LUCCa), a probabilistically-valid conformal prediction-based approach that accounts for epistemic uncertainty by calibrating the aleatoric uncertainty estimates provided by dynamics predictors. Unlike other CP methods [13, 20], we do not use a global calibration factor, as this tends to be overly conservative and uninformative. Instead our scaling factors are state and action dependent, thus accounting for changes in the approximate dynamics accuracy. LUCCa provably constructs calibrated regions for the first planning step, and, if the approximate dynamics are linear, also for all subsequent steps. We provide analysis showing the validity of our approach and present experiments where LUCCa is used to construct probabilistically-safe plans for a double-integrator under significant dynamics changes. Our contributions are the following:
@@ 5 LUCCa: Local Uncertainty Conformal Calibration
Given access to approximate dynamics $\tilde{f}$, and a finite calibration set of state transitions, we will show how to calibrate $\tilde{f}$’s estimates to be probabilistically valid for the epistemic uncertainty arising in inference. We can then chain prediction regions of the calibrated model in an MPC scheme to devise provably safe robot trajectories. Figure 1 showcases the overall method and Algorithm 1 details its steps.
@@ 5     $\hat{\mathcal{N}}_{\tau}\leftarrow\Sigma_{0}$ // Set starting state uncertainty, see Eq (11)
 // Propagate approximate dynamics with action sampled by MPPI
10 $\hat{\mathcal{N}}_{\tau+1}\leftarrow\tilde{f}(\hat{\mathcal{N}}_{\tau},u^{(sample)})$
@@ 5.2 Calibrating Uncertainty Estimates
Hence, scaling the approximate covariance matrix obtained from $\tilde{f}$ by the conformal scaling factor $\xi$ allows us to calibrate its uncertainty so that its $(1-\alpha)$ confidence region now contains at least $(1-\alpha)\%$ of the future states. This scaling factor acts as an interpretable measure of the model error of $\tilde{f}$. While the above enables the construction of a provably safe first prediction step, greater care must be taken when sequencing predictions over multiple steps. After the first prediction, the MVN passed as input to $\tilde{f}$ is not necessarily exchangeable with $\bar{D}_{cal}$ since it results from a different procedure, thus we do not provide a hard guarantee that the calibrated plans satisfy condition (2c) for arbitrary approximate dynamics $\tilde{f}$. However, we show in Appendix 0.A that, for linear $\tilde{f}$, our method provably leads to fully calibrated plans where every step satisfies condition (2c). We also empirically validate this property in §$\S 6$.
@@ 5.3 Local Calibration
In our implementation, we only pass the state mean and action as inputs to the decision tree, and thus do not separate between different initial uncertainties. Implicitly, the magnitude of the epistemic uncertainty must also be large enough when compared to the aleatoric uncertainty, to allow for informative feature space partitions. With this local procedure, our calibrated regions will not be undercovered for state-action pairs underrepresented in $D_{cal}$, might be more adaptive, and our planner has access to a scalar $\xi_{k}$ encapsulating the model’s accuracy. As when using SplitCP, the procedure above applies directly for one-step prediction no matter the form of $\tilde{f}$. However, we show also in Appendix 0.A that, for linear $\tilde{f}$, LUCCa provides local coverage guarantees over the entire planning horizon.
@@ 6 Experiments
We construct $D_{cal}$ by rolling out state-action pairs on a uniform grid for one step using $f$. To appropriately cover the space, we use a grid where $v_{x},v_{y}\in\{-1,-\frac{1}{2},0,\frac{1}{2},1\}$, $a_{x},a_{y}\in\{-0.8,-0.4,0,0.4,0.8\}$ and $p_{x},p_{y}\in{\{0,\frac{4.2}{15},\frac{8.4}{15},\ldots,4.2\}\cap\mathscr{C}}$. This ensures we appropriately cover the entire state-action space (with $\lvert\bar{D}_{cal}\rvert=53^{\prime}750$ for the Corridor map). In LOCART, we set the random split such that $\lvert\bar{D}_{cal,part}\rvert=\lvert\lceil 0.8\bar{D}_{cal}\rceil\rvert$, set the decision tree max depth to $13$ and the minimum number of samples required for a split to $40$. Empirically, if too small a ratio of calibration samples is used for feature space partition, the resulting $\xi_{k}$ plots might not (due to randomness in decision tree regressor and presence of aleatoric noise) properly match the true changes in model accuracy. For the parameters and $\lvert\bar{D}_{cal}\rvert$ above, calculating all the residuals and fitting the decision tree (creating the state-action space partitions and assigning a conformal scaling factor to each one) takes approximately $0.17$ secs on an Intel i9-12900K. Note this must only be performed once prior to planning (according to Algorithm 1). In Appendix 0.B, we show some $\xi_{k}$ obtained after fitting LOCART on each environment, and relate these scaling factor to the regions of high/low epistemic uncertainty.
The remainder of this section describes the two types of experiments we conducted. First, we sought to verify our claim that LUCCa produces a set that contains at least $(1-\alpha)$ percent of the outputs of the true dynamics for the first planning step. Additionally, since the double-integrator approximate dynamics $\tilde{f}$ are linear, the sets LUCCa produces for all subsequent planning steps should also contain at least $(1-\alpha)$ of the true dynamics outputs (as per Appendix 0.A). Second, we conducted experiments on an MPC controller that uses LUCCa to plan short-horizon trajectories to reach a goal. For both experiments, we compare to an uncalibrated baseline, i.e. directly using the uncertainty produced by the approximate dynamics, to show the importance of considering epistemic uncertainty. We use $\alpha=0.1$ for all experiments.
@@ 6.1 Empirical Coverage
To verify the central claim of this paper, that in the first planning step LUCCa outputs a set that contains at least $(1-\alpha)$ percent of the outputs of the true dynamics, we conducted an empirical coverage evaluation in the Corridor environment. Starting from two states (one in the white region $s_{0}=[3.0,3.6,0.5,0.0]^{\top}$ and one in the yellow $s_{0}=[2.1,1.5,0.0,0.5]^{\top}$), in the environment in Fig. 3, we applied a sequence of constant either accelerating or decelerating controls for the duration we use as our planning horizon. We accelerate and decelerate in the yellow region by ±$\pm 0.5$ m/s2, and in the white region by ±$\pm 0.8$ m/s2. We compare the ratio of $2500$ samples that fall within the predicted uncertainty region for each time step for LUCCa vs. the baseline in Fig. 4. For the baseline, the prediction region at each step can be obtained analytically by following the update rule $\hat{\Sigma}_{\tau+1}=A\hat{\Sigma}_{\tau}A^{\top}+Q$ and applying Lemma 2 to the resulting distribution. The figure shows that at least 90% of samples are within LUCCa’s prediction region in all cases, as expected, though LUCCa does appear to be more conservative than necessary, especially in later time steps. It also validates the extension of the $(1-\alpha)$ coverage guarantee to multi-step predictions for linear dynamics approximators. Conversely, the baseline, which does not consider epistemic uncertainty, significantly under-predicted the uncertainty for the state where the dynamics were inaccurate. For both states where $f=\tilde{f}$, we see that the baseline provides valid prediction regions (around $90\%$ of the samples are contained within the $90\%$ confidence hyperellipsoid) and LUCCa is slightly over-conservative. For the states where there’s significant model inaccuracy, LUCCa’s regions remain empirically valid for all planning steps while the baseline ones underestimate uncertainty in the true future states (with decaying validity over time).
@@ 6.2 Using LUCCa’s Predictions for Motion Planning
Figure 2 compares the trajectories produced by LUCCa and the baseline across the four environments. Videos of some executions can be found on our companion website um-arm-lab.github.io/lucca. The results in Table 1 suggest that using LUCCa’s uncertainty estimate improves the success rate in these scenarios by avoiding collision. However, LUCCa only guarantees that $(1-\alpha)$ percent of true states will be collision-free, so it does not provide a hard guarantee that the planned actions will result in collision-free states. We also note that the baseline did reach the goal faster when it didn’t collide. This is because its trace cost term is smaller and independent of the system’s current position, leading to greater velocities overall. On an Intel i9-12900K processor, the calibration process in LUCCa (passing MPPI sample through the decision tree to determine its leaf node and calibrating the approximate covariances with the corresponding scaling factor) was found to introduce a $0.3$ ms overhead on each planning step (averaged over $91000$ planning steps). While this represents a $32\%$ increase in the planning step duration, the magnitude of the additional computations is neglectable in practice (and could be further mitigated by GPU processing).
@@ 7 Conclusion
This paper presented LUCCa, a method based on conformal prediction that accounts for both aleatoric and epistemic uncertainty in the dynamics. The method uses a calibration set of data to locally estimate a scaling factor, which is applied to the output of the approximate dynamics predictor. Our experiments on a double-integrator system corroborate our analysis, showing that LUCCa outputs a set containing at least $(1-\alpha)$ percent of the results of the true dynamics for one-step predictions. Further, we prove and empirically observe that for linear approximate dynamics, LUCCa provides the same local coverage for the remaining planning steps. We also show that LUCCa’s predictions are effective for safely guiding an MPC method to reach a goal. {credits}

### v3

@@ Abstract
Whether learned, simulated, or analytical, approximations of a robot’s dynamics can be inaccurate when encountering novel environments. Many approaches have been proposed to quantify the aleatoric uncertainty of such methods, i.e. uncertainty resulting from stochasticity, however these estimates alone are not enough to properly estimate the uncertainty of a model in a novel environment, where the actual dynamics can change. Such changes can induce epistemic uncertainty, i.e. uncertainty due to a lack of information/data. Accounting for both epistemic and aleatoric dynamics uncertainty in a theoretically-grounded way remains an open problem. We introduce Local Uncertainty Conformal Calibration (LUCCa), a conformal prediction-based approach that calibrates the aleatoric uncertainty estimates provided by dynamics models to generate probabilistically-valid prediction regions of the system’s state. We account for both epistemic and aleatoric uncertainty non-asymptotically, without strong assumptions about the form of the true dynamics or how it changes. The calibration is performed locally in the state-action space, leading to uncertainty estimates that are useful for planning. We validate our method by constructing probabilistically-safe plans for a double-integrator under significant changes in dynamics. Companion website: https://um-arm-lab.github.io/lucca.
@@ 1 Introduction
This paper proposes Local Uncertainty Conformal Calibration (LUCCa), a probabilistically-valid conformal prediction-based approach that accounts for epistemic uncertainty by calibrating the aleatoric uncertainty estimates provided by dynamics predictors. Unlike other CP methods [14, 20], we do not use a global calibration factor, as this tends to be overly conservative and uninformative. Instead our scaling factors are state and action dependent, thus accounting for changes in the approximate dynamics accuracy. LUCCa provably constructs calibrated regions for the first planning step, and, if the approximate dynamics are linear and we can make certain assumptions about the controller, also for all subsequent steps. We provide analysis showing the validity of our approach and present experiments where LUCCa is used to construct probabilistically-safe plans for a double-integrator under significant dynamics changes. Our contributions are the following:
@@ 5 LUCCa: Local Uncertainty Conformal Calibration
Given access to approximate dynamics $\tilde{f}$, and a finite calibration set of state transitions, we will show how to calibrate $\tilde{f}$’s estimates to be probabilistically valid for the epistemic uncertainty arising in inference. We then chain prediction regions of the calibrated model in an MPC scheme which, under additional assumptions about the approximate dynamics and controller (see Appendix 0.A), devises provably safe robot trajectories. Figure 1 showcases the overall method and Algorithm 1 details its steps.
@@ 5     $\hat{\mathcal{N}}_{\tau}\leftarrow\Sigma_{0}$ // Set starting state uncertainty, see Eq (11)
10 $\hat{\mathcal{N}}_{\tau+1}\leftarrow\tilde{f}(\hat{\mathcal{N}}_{\tau},u^{(sample)})$ // Propagate $\hat{\mathcal{N}}_{\tau}$ with MPPI-sampled action
@@ 5.2 Calibrating Uncertainty Estimates
Hence, scaling the approximate covariance matrix obtained from $\tilde{f}$ by the conformal scaling factor $\xi$ allows us to calibrate its uncertainty so that its $(1-\alpha)$ confidence region now contains at least $(1-\alpha)\%$ of the future states. This scaling factor acts as an interpretable measure of the model error of $\tilde{f}$. While the above enables the construction of a provably safe first prediction step, greater care must be taken when sequencing predictions over multiple steps. After the first prediction, the MVN passed as input to $\tilde{f}$ is not necessarily exchangeable with $\bar{D}_{cal}$ since it results from a different procedure, thus we do not provide a hard guarantee that the calibrated plans satisfy condition (2c) for arbitrary approximate dynamics $\tilde{f}$. For guarantees about the multi-step case see Appendix 0.A. We empirically validate multi-step coverage in §$\S 6$.
@@ 5.3 Local Calibration
In our implementation, we only pass the state mean and action as inputs to the decision tree, and thus do not separate between different initial uncertainties. Implicitly, the magnitude of the epistemic uncertainty must also be large enough when compared to the aleatoric uncertainty, to allow for informative feature space partitions. With this local procedure, our calibrated regions will not be undercovered for state-action pairs underrepresented in $D_{cal}$, might be more adaptive, and our planner has access to a scalar $\xi_{k}$ encapsulating the model’s accuracy. As with SplitCP, the procedure above applies directly for one-step prediction no matter the form of $\tilde{f}$.
@@ 6 Experiments
We construct $D_{cal}$ by rolling out state-action pairs on a uniform grid for one step using $f$. To appropriately cover the space, we use a grid where $v_{x},v_{y}\in\{-1,-\frac{1}{2},0,\frac{1}{2},1\}$, $a_{x},a_{y}\in\{-0.8,-0.4,0,0.4,0.8\}$ and $p_{x},p_{y}\in{\{0,\frac{4.2}{15},\frac{8.4}{15},\ldots,4.2\}\cap\mathscr{C}}$. This ensures we appropriately cover the entire state-action space (with $\lvert\bar{D}_{cal}\rvert=53^{\prime}750$ for the Corridor map). In LOCART, we set the random split such that $\lvert\bar{D}_{cal,part}\rvert=\lvert\lceil 0.8\bar{D}_{cal}\rceil\rvert$, set the decision tree max depth to $13$ and the minimum number of samples required for a split to $40$. Empirically, if too small a ratio of calibration samples is used for feature space partition, the resulting $\xi_{k}$ plots might not (due to randomness in decision tree regressor and presence of aleatoric noise) properly match the true changes in model accuracy. For the parameters and $\lvert\bar{D}_{cal}\rvert$ above, calculating all the residuals and fitting the decision tree (creating the state-action space partitions and assigning a conformal scaling factor to each one) takes approximately $0.17$ secs on an Intel i9-12900K. Note this must only be performed once prior to planning (according to Algorithm 1). In Appendix 0.B, we show some $\xi_{k}$ obtained after fitting LOCART on each environment, and relate these scaling factors to the regions of high/low epistemic uncertainty.
The remainder of this section describes the two types of experiments we conducted. First, we sought to verify our claim that LUCCa produces a set that contains at least $(1-\alpha)$ percent of the outputs of the true dynamics for the first planning step. Additionally, although the presented system with bounded control inputs and $\dim(\mathcal{S})>\dim(\mathcal{U})$ is more complex than those for which we provably extended LUCCa’s safety guarantees to the multi-step case (Appendix 0.A), we wanted to test whether LUCCa could still provide $(1-\alpha)$ percent safety along each step of its plan. Second, we conducted experiments on an MPC controller that uses LUCCa to plan short-horizon trajectories to reach a goal. For both experiments, we compare to an uncalibrated baseline, i.e. directly using the uncertainty produced by the approximate dynamics, to show the importance of considering epistemic uncertainty. We use $\alpha=0.1$ for all experiments.
@@ 6.1 Empirical Coverage
To verify the central claim of this paper, that in the first planning step LUCCa outputs a set that contains at least $(1-\alpha)$ percent of the outputs of the true dynamics, we conducted an empirical coverage evaluation in the Corridor environment. Starting from two states (one in the white region $s_{0}=[3.0,3.6,0.5,0.0]^{\top}$ and one in the yellow $s_{0}=[2.1,1.5,0.0,0.5]^{\top}$), in the environment in Fig. 3, we applied a sequence of constant either accelerating or decelerating controls for the duration we use as our planning horizon. We accelerate and decelerate in the yellow region by ±$\pm 0.5$ m/s2, and in the white region by ±$\pm 0.8$ m/s2. We compare the ratio of $2500$ samples that fall within the predicted uncertainty region for each time step for LUCCa vs. the baseline in Fig. 4. For the baseline, the prediction region at each step can be obtained analytically by following the update rule $\hat{\Sigma}_{\tau+1}=A\hat{\Sigma}_{\tau}A^{\top}+Q$ and applying Lemma 2 to the resulting distribution. The figure shows that at least 90% of samples are within LUCCa’s first-step prediction regions in all cases, as expected, though LUCCa does appear to be more conservative than necessary (especially in later time steps). It also validates that, despite the tested system being more complex than those for which we proved multi-step safety, LUCCa appears to also provide multi-step local coverage for linear approximate dynamics. Conversely, the baseline, which does not consider epistemic uncertainty, significantly under-predicted the uncertainty for the state where the dynamics were inaccurate. For both states where $f=\tilde{f}$, we see that the baseline provides valid prediction regions (around $90\%$ of the samples are contained within the $90\%$ confidence hyperellipsoid) and LUCCa is slightly over-conservative. For the states where there’s significant model inaccuracy, LUCCa’s regions remain empirically valid for all planning steps while the baseline ones underestimate uncertainty in the true future states (with decaying validity over time).
@@ 6.2 Using LUCCa’s Predictions for Motion Planning
Figure 2 compares the trajectories produced by LUCCa and the baseline across the four environments. Videos of some executions can be found on our companion website https://um-arm-lab.github.io/lucca. The results in Table 1 suggest that using LUCCa’s uncertainty estimate improves the success rate in these scenarios by avoiding collision. However, LUCCa only guarantees that $(1-\alpha)$ percent of true states will be collision-free, so it does not provide a hard guarantee that the planned actions will result in collision-free states. We also note that the baseline did reach the goal faster when it didn’t collide. This is because its trace cost term is smaller and independent of the system’s current position, leading to greater velocities overall. On an Intel i9-12900K processor, the calibration process in LUCCa (passing MPPI sample through the decision tree to determine its leaf node and calibrating the approximate covariances with the corresponding scaling factor) was found to introduce a $0.3$ ms overhead on each planning step (averaged over $91000$ planning steps). While this represents a $32\%$ increase in the planning step duration, the magnitude of the additional computations is neglectable in practice (and could be further mitigated by GPU processing).
@@ 7 Conclusion
This paper presented LUCCa, a method based on conformal prediction that accounts for both aleatoric and epistemic uncertainty in the dynamics. The method uses a calibration set of data to locally estimate a scaling factor, which is applied to the output of the approximate dynamics predictor. Our experiments on a double-integrator system corroborate our analysis, showing that LUCCa outputs a set containing at least $(1-\alpha)$ percent of the results of the true dynamics for one-step predictions. Further, we prove and empirically observe that for linear approximate dynamics, under additional controller assumptions, LUCCa can also provide local coverage for the remaining planning steps. We also show that LUCCa’s predictions are effective for safely guiding an MPC method to reach a goal. {credits}